OpenStudy (anonymous):

Two trucks started toward each other at the same time from towns 500 km apart. One truck traveled at a rate of 65 km per hour while the other traveled at 60 km per hour. After how many hours did they meet? If x represents the number of hours, which of the following equations could be used to solve the problem? A. 65x = 60x B. 65x - 60x = 500 C. 65x + 60x = 500

OpenStudy (anonymous):

The answer would be \[65x+10x=500\] The way I explain it is, you can think of one of the trucks as standing still and the other one approaching at the sum of their speeds That gives you \[{65+60}\times x =500\] for the moving truck. Now convert it to \[65x+10x=500\]